OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer. 10 - 2 * log x = 13

OpenStudy (anonymous):

...

OpenStudy (anonymous):

\[10-2\log(x)=13\] \[-2\log(x)=3\] \[\log(x)=-\frac{3}{2}\] \[x=10^{-\frac{3}{2}}\]

OpenStudy (anonymous):

2*log(x)=-3 log(x)=-1.5 x=10^-1.5 so the answer is 0.0316228